# Non-Gaussian state teleportation with a nonlinear feedforward

### v3

@@ 1 Introduction
Entangled Gaussian cluster states offer a rich multimode structure. Accompanied by single photon measurements, their simulation complexity exceeds the abilities of classical computation [12, 5]. Their capacity for massive multi-mode multiplexing [32, 48, 22, 28] is the main advantage of measurement-induced quantum computing with bosonic modes of traveling light. In this protocol, measurements of the individual modes enable the quantum information to propagate along the cluster and evolve in the desired manner [29, 28, 47]. The choice of measurement is what determines whether the state is simply transmitted by quantum teleportation [2, 8, 35, 47, 41] or transformed [37, 7, 31, 13]. To reach the full advantages of the quantum computation, the Gaussian entangled states need to be further supported by deterministic non-Gaussian operations [25, 11, 26, 5]. These challenging operations can be implemented in a measurement-induced fashion, [31, 18] leaving it with two distinct tasks: experimental generation of the non-Gaussian states, and successful propagation of the non-Gaussian states through a Gaussian circuit.
@@
In quantum optics with traveling light, non-Gaussian quantum states have been generated in a wide range of experimental configurations [49, 44, 33, 50, 34, 21]. Today, photon-number-resolving detectors, together with Gaussian operations such as parametric down-conversion and coherent classical displacements, are feasibly used to probabilistically generate non-Gaussian states with varied nonclassical and non-Gaussian properties [19, 42, 1, 27, 43, 20]. One of these important properties is the nonlinear squeezing that can be used as a resource for deterministic preparation of cubic nonlinearity [31, 36] and is a specific example of a purely continuous non-Gaussian resource [17, 19, 45]. The basic principle of the state preparation, closely aligned to the Gaussian boson sampling problem [12], is still actively investigated to prepare more and more complex quantum states [39]. The successful generation of non-Gaussian states led to their active use in tests of their propagation through Gaussian circuits, resulting in demonstration of key protocols, such as teleportation [15, 30, 9, 40, 3, 23, 51], decoherence prevention [16], or nonlinear measurement [36]. One thing shared by all of these experimental tests is that the non-Gaussian features of the employed states are deteriorated in the process, diminishing their power. Therefore, novel ways of processing that help preserve the non-Gaussianity will be needed while moving forward.
@@
In this paper we are going to show how the canonical teleportation protocol along a single node of an entangled cluster can be enhanced by a classical non-linear feedforward to better transfer a specific class of non-Gaussian quantum states. We focus on quantum states with cubic nonlinear squeezing which is a resource for measurement-based quantum computation with continuous variables. To this end we are going to consider the problem of teleporting these states, substantially different than Fock states, qubit states and cat states, and compare the canonical teleportation to the teleportation with enhanced feedforward operation. In the field of measurement-induced operations, the teleportation of a state with nonlinear squeezing enriched by nonlinear feedforward can be seen as an approximation of the cubic nonlinear phase measurement [36]. The teleported state serves as an ancillary state and together with the homodyne measurements and feedforward they construct the approximated measurement which applies to the first mode of the two-mode Gaussian state. Therefore, our analyses will also test the projection powers of the approximate measurement as a state preparation. We are going to compare this approximation to the ultimate ideal theoretical bound provided by the ideal cubic measurement.
@@ 2.1 Deterministic regime
The teleportation protocol [8, 35] is the basic building block of the cluster state computing. The usual involved measurement is the homodyne detection with a linear feedforward that provides linear dependence of operations performed on the outcoming state on measurements results. It is possible to consider also generalized measurement, that enables not only transfer of quantum states but also their transformation. Such a measurement can include adjustable nonlinear feedforward that utilizes the measurement results in some more complex function.
@@
|$\ket{\chi}=C(\chi)S(r)\ket{0},$ (2)
where |$\ket{p=0}$ is an eigenstate of the $p$ quadrature with eigenvalue 0. In this case, the teleportation can be seen as similar to remote state preparation, which could be best achieved by measuring one of the entangled modes in a correct basis. The projection on a cubic state can be performed by the ideal cubic measurement [36], illustrated in Fig. 1c. POVM element of the measurement can be expressed as
$\Pi=C(\chi)\ket{p=m}\bra{p=m}C(\chi)^{\dagger},$ (3)
@@
is the cubic nonlinearity and |$\ket{p=m}$ is an eigenstate of the $p$ quadrature with eigenvalue $m$. In the ideal scenario, the measurement can be understood as measurement of nonlinear combination of quadrature operators
@@
that projects on its eigenstates. Such states have infinite energy, similar to position eigenstates, which are projected on by homodyne measurement [11, 10]. The measurement can be decomposed into the cubic nonlinearity and p-homodyne measurement. These operations can be sequentially applied to one part of the two-mode cluster state as is shown in Fig. 1c. It is necessary to utilize the measurement result $m$ in a Gaussian feedforward in order to obtain a non-Gaussian state. If the measurement outcome is disregarded, the outcoming state is quantum Gaussian, i.e. it is a mixture of Gaussian states. In Heisenberg picture, the overall transformation yields
@@ 2.3 Nonlinear squeezing
The original role of non-Gaussian ancillary state is to suppress a noise term that appears at the outcome of the measurement-induced cubic phase gate [31, 19]. This term has the same form as the operator (5) that is measured by the nonlinear measurement; the real parameter $z$ is called cubicity. Inspired by insufficiency of Gaussian states for the quantum computing advantage with CV, the nonlinear squeezing is defined as a ratio of variance of the operator $O$ compared to the minimal variance that can be achieved in a Gaussian state [17, 45, 19]
@@
Experimentally feasible approximations of the finitely nonlinearly squeezed states can be prepared on a finite Fock subspace [49, 14]. The approximations can be written as
|$\ket{\phi_{N}}=\sum_{n=0}^{N}c_{n}\ket{n},$ (19)
where the coefficients can be found in an efficient way by solving an eigenvalue problem and two variable optimization [31]. Note that Fock states and states with nonlinear squeezing have different symmetries. Unlike the nonlinearly squeezed states that are symmetric in $x$ only, Wigner functions of Fock states are symmetric both in $x$ and $p$. This difference hints that an approximation of the cubic state must be a superposition of Fock states. The experimentally feasible quantum state exhibiting nonlinear squeezing and simultaneously the lowest order approximation is a superposition of the first two Fock states
|$\ket{\phi_{2}}=u\ket{0}+i\sqrt{1-u^{2}}\ket{1},$ (20)
where $u$ is real and bounded by $10<u<$ 1 [19]. The used parameterization ensures symmetry under transformation $x\rightarrow-x$ which holds also for (5). One step further leads to ancillary state approximation
@@
where the $c_{i}$ are real coefficients. The states can be found as eigenstates of the operator $O$ restricted to an $N$ dimensional subspace of the Fock space. They provide the best nonlinear squeezing available on such a subspace [31].
To show the positive effect of nonlinear feedforward, let us compare the results obtained by the regular and nonlinear teleportation protocols. At the same time, we compare the schemes equipped with identical resources of non-Gaussianity [38] and Gaussian squeezing [4]. The approximations can be classified by the faithful hierarchy of $N$-photon non-Gaussianity [21] or stellar rank [6]. A superposition (19) with a fixed highest Fock state $N$ has $N$-photon non-Gaussianity. When teleporting a state with some given $N$-photon non-Gaussianity and transferring a state from the same level of the hierarchy by nonlinear teleportation, both schemes posses the same $N$-photon non-Gaussian state and any improvement in performance thus originates from the properties of the scheme. Another resource present in the schemes is Gaussian squeezing [4]. Throughout our analysis we limit the available Gaussian squeezing while comparing the performances. At first in preparation of the finitely squeezed state (18) and as well of the cluster state. We consider a general two-mode Gaussian cluster state, whose parameters are optimized for each situation and each scheme. The optimization is done with help of Python library scipy optimize and repeated with randomly chosen initial conditions. Details can be found in Appendix C. To remain close to the experimental reality, we consider an upperbound on Gaussian squeezing available. The optimum is then sought in a range given by these bounds, e.g. $\pm$9 dB. Overall, by limiting the resources we can perform a genuine comparison of schemes equipped with the same level of $N$-photon non-Gaussianity and limit on maximal Gaussian squeezing.
The parameterization of the two-mode Gaussian state via the covariance matrix can be found in Appendix B and is equivalent to two thermal states transformed by single mode Gaussian operations and a beam splitter as shown in Fig. 1b. The thermal noise is taken as an input parameter and by its change, we can analyze the resilience of discussed effects. It is described by parameter $n$ representing the multiples of vacuum fluctuations added to an initially pure vacuum state. Thus, for a pure state the parameter yields $n=0$ , while $n>0$ represents some amount of thermal noise [46].
@@ 3.2 Probabilistic regime
Nonlinear squeezing shown in Fig. 6 is visibly enhanced by nonlinear feedforward and enables its detection even for probabilities for which the regular teleportation fails to transfer any. We also introduced $25\%$ losses to the ancillary state, that was present in its recent experimental preparation. The results are shown in dashed lines and again show advantage of the nonlinear teleportation with nonlinear feedforward. The optimal strength of the nonlinear feedforward is higher than for the ideal case, yielding $\chi=-0.250$ for 6dB cluster state and $\chi=-0.151$ for 8dB cluster state. A similar effect was recognized in decoherence of the nonlinear squeezing [17], where the native cubicity of the state increases with higher losses. In the presence of losses, the performances of teleportations through cluster states with 6 and 8 dB of Gaussian squeezing does not converge even for postselection on the best cases with vanishing probability.
Figure 6: Nonlinear squeezing (16) produced by unity gain regular/nonlinear teleportation of an eigenstate of the operator $O$ (5) conditioned on measurement results of the homodyne measurement with best results aggregated up to some probability and evaluated at native cubicity of the state before teleportation. The nonlinear teleportation and regular teleportation was simulated with two mode squeezed vacuum of 6 dB and 8 dB with (dashed lines) and without losses on the teleported state. Here it is assumed $25\%$ losses [19]. In all cases the nonlinear squeezing firstly vanishes (the plot line crosses the dashed line of boundary 1) for regular teleportation.
@@ 4 Conclusion
We showed an improvement in the teleported nonlinear squeezing by involving currently experimentally tested nonlinear feedforward [36]. In a deterministic regime, the positive effect can be seen on the theoretical level of teleportation of the ideal cubic state. However in a probabilistic regime, when best results are postselected up to some probability, the improvement is observable even with current experimental resources in comparison to unity gain teleportation. Processing the measurement data during teleportation via classical nonlinear feedforward decreases the added amount of noise which can be seen by improved nonlinear squeezing at the output. The cubic nonlinear squeezing is a resource for measurement-based quantum computation [19, 36] and improving its amount that is transferred via teleportation increases the potential to chain and iterate computational steps as required by cluster computing.
The cubic nonlinearity itself is not currently experimentally available in travelling light, however a superposition of Fock states with nonlinear squeezing has already been prepared [19] as an alternative resource substituting the nonlinearity. Secondly, we show that in probabilistic regime the nonlinear teleportation equipped with nonlinear feedforward and currently experimentaly accessible state outperforms the unity gain linear teleportation through 6dB cluster state as is shown in Fig. 6. The advantage is thus in principle observable even with currently available nonlinearly squeezed state and appears by adjusting the classical part of the scheme consisting of nonlinear feedforward combined with post-selection. The enhancement remains observable even in the presence of losses that accompanied the experimental preparation [19]. Our work is one of the first steps required to extend the presently Gaussian theory of cluster states into its final destination - non-Gaussian domain.
@@ Acknowledgement
VK and PM acknowledge grant 22-08772S and RF grant 21-13491X of Czech Science Foundation. VK, PM and RF also acknolwedge Horizon Europe Research and Innovation Actions under Grant Agreement no. 101080173 (CLUSTEC). P.M. and R.F. acknowledge project CZ.02.01.010022 0080004649 (QUEENTEC) of EU and the Czech Ministry of Education, Youth and Sport. M.W. acknowledges funding from the Plan France 2030 through the project ANR-23-PETQ-0013. VK acknowledges project IGA-PrF-2024-008.
@@ Appendix A Model in Heisenberg picture
The Gaussianity of the cluster state can be used for expressing the whole expression in terms appearing in covariance matrix of the modes A and B. For the more complex moments involving more than two quadratures, we use the generalized Stein’s lemma for multivariate Gaussian distributions of a random variable $\vec{X}$ with mean values $\vec{\mu}$ [24]
@@ Appendix C Optimization of the cluster state
We performed numerical optimization with scipy.optimize minimize modul with 400-460 randomly chosen initial states with covariance matrix (47). The initial conditions were chosen within a range given by the upper bound on the maximal Gaussian squeezing, an interval ±$\pm 10$ for mean values of $x$ quadratures, cubicities $\chi$ and $z$ and gains in an interval ±$\pm 2$ and arbitrary rotations.

### v4

@@ 1 Introduction
Entangled Gaussian cluster states offer a rich multimode structure. Accompanied by single photon measurements, their simulation complexity exceeds the abilities of classical computation [GBS, Chabaud2023]. Their capacity for massive multi-mode multiplexing [OBrien2007, Yokohama2013, Larsen, Menicucci] is the main advantage of measurement-induced quantum computing with bosonic modes of traveling light. In this protocol, measurements of the individual modes enable the quantum information to propagate along the cluster and evolve in the desired manner [ClusterCV, Menicucci, Walshe2020]. The choice of measurement is what determines whether the state is simply transmitted by quantum teleportation [Ban2002, Furusawa1998, Pirandola2015, Walshe2020, Takeoka] or transformed [Sefi2019, Filip2005, Miyata2016, Hillmann2022]. To reach the full advantages of the quantum computation, the Gaussian entangled states need to be further supported by deterministic non-Gaussian operations [Lloyd1999, Gottesman2001, Mari2012, Chabaud2023]. These challenging operations can be implemented in a measurement-induced fashion, [Miyata2016, Konno2021b] leaving it with two distinct tasks: experimental generation of the non-Gaussian states, and successful propagation of the non-Gaussian states through a Gaussian circuit.
@@
In quantum optics with traveling light, non-Gaussian quantum states have been generated in a wide range of experimental configurations [Yukawa2013b, URen2004, Ogawa2016, Zavatta2004, Ourjoumtsev2006, LachmanG19]. Today, photon-number-resolving detectors, together with Gaussian operations such as parametric down-conversion and coherent classical displacements, are feasibly used to probabilistically generate non-Gaussian states with varied nonclassical and non-Gaussian properties [Konno2021, Takeoka2011, Asavanant17, Melalkia22, Tiedau, Konno2024]. One of these important properties is the nonlinear squeezing that can be used as a resource for deterministic preparation of cubic nonlinearity [Miyata2016, sakaguchi] and is a specific example of a purely continuous non-Gaussian resource [Kala22, Konno2021, Brauer2021]. The basic principle of the state preparation, closely aligned to the Gaussian boson sampling problem [GBS], is still actively investigated to prepare more and more complex quantum states [Takase2023]. The successful generation of non-Gaussian states led to their active use in tests of their propagation through Gaussian circuits, resulting in demonstration of key protocols, such as teleportation [Ide, Mista, Fuwa, Takeda, Benichi, Lee, Zhao2023], decoherence prevention [Jeannic2018], or nonlinear measurement [sakaguchi]. One thing shared by all of these experimental tests is that the non-Gaussian features of the employed states are deteriorated in the process, diminishing their power. Therefore, novel ways of processing that help preserve the non-Gaussianity will be needed while moving forward.
@@
In this paper we are going to show how the canonical teleportation protocol along a single node of an entangled cluster can be enhanced by a classical non-linear feedforward to better transfer a specific class of non-Gaussian quantum states. We focus on quantum states with cubic nonlinear squeezing which is a resource for measurement-based quantum computation with continuous variables. To this end we are going to consider the problem of teleporting these states, substantially different than Fock states, qubit states and cat states, and compare the canonical teleportation to the teleportation with enhanced feedforward operation. In the field of measurement-induced operations, the teleportation of a state with nonlinear squeezing enriched by nonlinear feedforward can be seen as an approximation of the cubic nonlinear phase measurement [sakaguchi]. The teleported state serves as an ancillary state and together with the homodyne measurements and feedforward they construct the approximated measurement which applies to the first mode of the two-mode Gaussian state. Therefore, our analyses will also test the projection powers of the approximate measurement as a state preparation. We are going to compare this approximation to the ultimate ideal theoretical bound provided by the ideal cubic measurement.
@@ 2.1 Deterministic regime
The teleportation protocol [Furusawa1998, Pirandola2015] is the basic building block of the cluster state computing. The usual involved measurement is the homodyne detection with a linear feedforward that provides linear dependence of operations performed on the outcoming state on measurements results. It is possible to consider also generalized measurement, that enables not only transfer of quantum states but also their transformation. Such a measurement can include adjustable nonlinear feedforward that utilizes the measurement results in some more complex function.
@@
[…]
@@
[…]
@@
that projects on its eigenstates. Such states have infinite energy, similar to position eigenstates, which are projected on by homodyne measurement [Gottesman2001, Ghose2007]. The measurement can be decomposed into the cubic nonlinearity and p-homodyne measurement. These operations can be sequentially applied to one part of the two-mode cluster state as is shown in Fig. 1c. It is necessary to utilize the measurement result $m$ in a Gaussian feedforward in order to obtain a non-Gaussian state. If the measurement outcome is disregarded, the outcoming state is quantum Gaussian, i.e. it is a mixture of Gaussian states. In Heisenberg picture, the overall transformation yields
@@ 2.3 Nonlinear squeezing
The original role of non-Gaussian ancillary state is to suppress a noise term that appears at the outcome of the measurement-induced cubic phase gate [Miyata2016, Konno2021]. This term has the same form as the operator (5) that is measured by the nonlinear measurement; the real parameter $z$ is called cubicity. Inspired by insufficiency of Gaussian states for the quantum computing advantage with CV, the nonlinear squeezing is defined as a ratio of variance of the operator $O$ compared to the minimal variance that can be achieved in a Gaussian state [Kala22, Brauer2021, Konno2021]
@@
Experimentally feasible approximations of the finitely nonlinearly squeezed states can be prepared on a finite Fock subspace [Yukawa2013b, Huang]. The approximations can be written as
@@
where the coefficients can be found in an efficient way by solving an eigenvalue problem and two variable optimization [Miyata2016]. Note that Fock states and states with nonlinear squeezing have different symmetries. Unlike the nonlinearly squeezed states that are symmetric in $x$ only, Wigner functions of Fock states are symmetric both in $x$ and $p$. This difference hints that an approximation of the cubic state must be a superposition of Fock states. The experimentally feasible quantum state exhibiting nonlinear squeezing and simultaneously the lowest order approximation is a superposition of the first two Fock states
[…]
where $u$ is real and bounded by $0<u<1$ [Konno2021]. The used parameterization ensures symmetry under transformation $x\rightarrow-x$ which holds also for (5). One step further leads to ancillary state approximation
[…]
where the $c_{i}$ are real coefficients. The states can be found as eigenstates of the operator $O$ restricted to an $N$ dimensional subspace of the Fock space. They provide the best nonlinear squeezing available on such a subspace [Miyata2016].
To show the positive effect of nonlinear feedforward, let us compare the results obtained by the regular and nonlinear teleportation protocols. At the same time, we compare the schemes equipped with identical resources of non-Gaussianity [Takagi] and Gaussian squeezing [Braunstein2005b]. The approximations can be classified by the faithful hierarchy of $N$-photon non-Gaussianity [LachmanG19] or stellar rank [chabaud2020]. A superposition (19) with a fixed highest Fock state $N$ has $N$-photon non-Gaussianity. When teleporting a state with some given $N$-photon non-Gaussianity and transferring a state from the same level of the hierarchy by nonlinear teleportation, both schemes posses the same $N$-photon non-Gaussian state and any improvement in performance thus originates from the properties of the scheme. Another resource present in the schemes is Gaussian squeezing [Braunstein2005b]. Throughout our analysis we limit the available Gaussian squeezing while comparing the performances. At first in preparation of the finitely squeezed state (18) and as well of the cluster state. We consider a general two-mode Gaussian cluster state, whose parameters are optimized for each situation and each scheme. The optimization is done with help of Python library scipy optimize and repeated with randomly chosen initial conditions. Details can be found in Appendix C. To remain close to the experimental reality, we consider an upperbound on Gaussian squeezing available. The optimum is then sought in a range given by these bounds, e.g. $\pm$9 dB. Overall, by limiting the resources we can perform a genuine comparison of schemes equipped with the same level of $N$-photon non-Gaussianity and limit on maximal Gaussian squeezing.
The parameterization of the two-mode Gaussian state via the covariance matrix can be found in Appendix B and is equivalent to two thermal states transformed by single mode Gaussian operations and a beam splitter as shown in Fig. 1b. The thermal noise is taken as an input parameter and by its change, we can analyze the resilience of discussed effects. It is described by parameter $n$ representing the multiples of vacuum fluctuations added to an initially pure vacuum state. Thus, for a pure state the parameter yields $n=0$ , while $n>0$ represents some amount of thermal noise [Mattia2020].
@@ 3.2 Probabilistic regime
Nonlinear squeezing shown in Fig. 6 is visibly enhanced by nonlinear feedforward and enables its detection even for probabilities for which the regular teleportation fails to transfer any. We also introduced $25\%$ losses to the ancillary state, that was present in its recent experimental preparation. The results are shown in dashed lines and again show advantage of the nonlinear teleportation with nonlinear feedforward. The optimal strength of the nonlinear feedforward is higher than for the ideal case, yielding $\chi=-0.250$ for 6dB cluster state and $\chi=-0.151$ for 8dB cluster state. A similar effect was recognized in decoherence of the nonlinear squeezing [Kala22], where the native cubicity of the state increases with higher losses. In the presence of losses, the performances of teleportations through cluster states with 6 and 8 dB of Gaussian squeezing does not converge even for postselection on the best cases with vanishing probability.
Figure 6: Nonlinear squeezing (16) produced by unity gain regular/nonlinear teleportation of an eigenstate of the operator $O$ (5) conditioned on measurement results of the homodyne measurement with best results aggregated up to some probability and evaluated at native cubicity of the state before teleportation. The nonlinear teleportation and regular teleportation was simulated with two mode squeezed vacuum of 6 dB and 8 dB with (dashed lines) and without losses on the teleported state. Here it is assumed $25\%$ losses [Konno2021]. In all cases the nonlinear squeezing firstly vanishes (the plot line crosses the dashed line of boundary 1) for regular teleportation.
@@ 4 Conclusion
We showed an improvement in the teleported nonlinear squeezing by involving currently experimentally tested nonlinear feedforward [sakaguchi]. In a deterministic regime, the positive effect can be seen on the theoretical level of teleportation of the ideal cubic state. However in a probabilistic regime, when best results are postselected up to some probability, the improvement is observable even with current experimental resources in comparison to unity gain teleportation. Processing the measurement data during teleportation via classical nonlinear feedforward decreases the added amount of noise which can be seen by improved nonlinear squeezing at the output. The cubic nonlinear squeezing is a resource for measurement-based quantum computation [Konno2021, sakaguchi] and improving its amount that is transferred via teleportation increases the potential to chain and iterate computational steps as required by cluster computing.
The cubic nonlinearity itself is not currently experimentally available in travelling light, however a superposition of Fock states with nonlinear squeezing has already been prepared [Konno2021] as an alternative resource substituting the nonlinearity. Secondly, we show that in probabilistic regime the nonlinear teleportation equipped with nonlinear feedforward and currently experimentaly accessible state outperforms the unity gain linear teleportation through 6dB cluster state as is shown in Fig. 6. The advantage is thus in principle observable even with currently available nonlinearly squeezed state and appears by adjusting the classical part of the scheme consisting of nonlinear feedforward combined with post-selection. The enhancement remains observable even in the presence of losses that accompanied the experimental preparation [Konno2021]. Our work is one of the first steps required to extend the presently Gaussian theory of cluster states into its final destination - non-Gaussian domain.
@@ Acknowledgement
VK and PM acknowledge grant 22-08772S and RF grant 21-13491X of Czech Science Foundation. VK acknowledges the Quantera project CLUSSTAR (8C24003) of MEYS, Czech Republic. Project CLUSSTAR has received funding from the European Union’s Horizon 2020 Research and Innovation Programme under Grant Agreements No. 731473 and No. 101017733 (QuantERA). PM acknowledges a grant from the Programme Johannes Amos Comenius under the Ministry of Education, Youth and Sports of the Czech Republic reg. no. CZ.02.01.01/00/22_008/0004649. VK, PM and RF also acknolwedge Horizon Europe Research and Innovation Actions under Grant Agreement no. 101080173 (CLUSTEC). P.M. and R.F. acknowledge project CZ.02.01.010022 0080004649 (QUEENTEC) of EU and the Czech Ministry of Education, Youth and Sport. M.W. acknowledges funding from the Plan France 2030 through the project ANR-23-PETQ-0013. VK acknowledges project IGA-PrF-2024-008.
@@ Appendix A Model in Heisenberg picture
The Gaussianity of the cluster state can be used for expressing the whole expression in terms appearing in covariance matrix of the modes A and B. For the more complex moments involving more than two quadratures, we use the generalized Stein’s lemma for multivariate Gaussian distributions of a random variable $\vec{X}$ with mean values $\vec{\mu}$ [Liu1994]
@@ Appendix C Optimization of the cluster state
We performed numerical optimization with scipy.optimize minimize modul with 400-460 randomly chosen initial states with covariance matrix (47). The initial conditions were chosen within a range given by the upper bound on the maximal Gaussian squeezing, an interval $\pm 10$ for mean values of $x$ quadratures, cubicities $\chi$ and $z$ and gains in an interval $\pm 2$ and arbitrary rotations.